The total area of a cube is 216 cm2. Find the length of an edge.
Answer by mananth(16946)


The total area of a cube is 216 cm2. Find the length of an edge.
Total area = 6 l^2 l =length of side/edge
6l^2 =216
l^2 = 216/6 = 36
l^2 = 36
l =6 cm